In an exam 70% passed in English 65% passed in Maths while 27% failed in both.If 248 candidates  passed in both.Find the total number of candidates

Grade:12

1 Answers

Priyansh Bajaj AskiitiansExpert-IITD
13 years ago

Dear Vivek,

Solution:-Let the total no. of candidates are 'n'.

Then, no. of candidates passed in English (E) = 0.7n,

        no. of candidates passed in Maths (M) = 0.65n, and

        no. of candidates neither passed in English nor in Maths (E‾ ∩ M‾) = 0.27n.

[Here, E‾ and  M‾ are the complements of E and M respectively.]

        no. of candidates passed in both (E ∩ M) = 248

We know, (E U M) = n - (E‾ ∩ M‾) [Very standard formula]

and (E U M) = (E) + (M) - (E ∩ M) [Another standard formula]

These two formulas will give-

n - (E‾ ∩ M‾) = (E) + (M) - (E ∩ M)

n - 0.27n = 0.7n+0.65n - 248

n = 400 (Ans)


Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We are all IITians and here to help you in your IIT JEE preparation.


All the best Vivek!!!


Regards,

Askiitians Experts

Priyansh Bajaj